# Natural Metric-Affine Inflation

### v1

@@ 5 Conclusions
We have focused on the low-energy (two-derivative) metric-affine inflationary theory containing only the graviton and the PNGB inflaton in the particle content. In particular, the connection does not carry additional degrees of freedom. This theory, whose action is given in (2.5), features, besides the PNGB potential, non-minimal couplings between the PNGB and the two linear-in-curvature invariants that one can construct in metric-affine gravity: ${\cal R}$ and $\tilde{\cal R}$, defined in (2.7). We have integrated out the distorsion to find an explicit metric theory, with action given in (2.11), where the connection is given by the LC formula. Moreover, in Appendix A we have found an explicit UV completion for the cosine potential and both non-minimal couplings in (2.6) and (2.20), respectively.
@@
Finally, we have performed an analysis of the inflationary predictions of this theory for the potential and non-minimal couplings in (2.6) and (2.20), respectively, because, as shown, they admit an explicit UV completion. For these models, we have found regions of the parameter space where the inflationary predictions agree with the most recent observations performed by the Planck, BICEP and Keck collaborations at the $2\sigma$ level. We have found that in order to enter the $1\sigma$ region it is necessary (and sufficient) to have a finite value of the Barbero-Immirzi parameter (a sizable $|\beta_{0}|$) and a sizable $\tilde{\xi}$ (with sign opposite to the Barbero-Immirzi parameter). This is precisely the region of parameter space for which the potential of $\chi$ develops a plateau. It is indeed the presence of this plateau that renders the theory observationally viable.
@@ Acknowledgments
The work of A.R. was supported by the Estonian Research Council grants PRG1055, RVTT3, RVTT7 and the CoE program TK202 “Fundamental Universe”. The work of A.S. was also supported by the grant DyConn from the University of Rome Tor Vergata.
@@ Appendix A A microscopic origin of natural metric-affine inflation
This relation should also be approximately true in this variant of QCD with $f$ taken at the Planck scale as long as the elements of $M_{q}$ are much smaller than $f$.

### v2

@@ 5 Conclusions
We have focused on the low-energy (two-derivative) metric-affine inflationary theory containing only the graviton and the PNGB inflaton in the particle content. In particular, the connection does not carry additional degrees of freedom. This theory, whose action is given in [^5] It is worth mentioning that (2.5) is also the low-energy metric-affine inflationary theory for any single-field scenario with generic potential $V$, which is not necessarily the natural-inflation potential in (2.6). (2.5), features, besides the PNGB potential, non-minimal couplings between the PNGB and the two linear-in-curvature invariants that one can construct in metric-affine gravity: ${\cal R}$ and $\tilde{\cal R}$, defined in (2.7). We have integrated out the distorsion to find an explicit metric theory, with action given in (2.11), where the connection is given by the LC formula. Moreover, in Appendix A we have found an explicit UV completion for the cosine potential and both non-minimal couplings in (2.6) and (2.20), respectively: by UV completion we mean that these functions emerge from the microscopic dynamics of a quantum field theory that is renormalizable on Minkowski spacetime. The UV completion of Einstein gravity is beyond the scope of the present paper; perhaps it can be achieved in string theory or in asymptotically free/safe theories featuring higher-derivative terms.
@@
Finally, we have performed an analysis of the inflationary predictions of this theory for the potential and non-minimal couplings in (2.6) and (2.20), respectively, because, as shown, they admit an explicit UV completion. For these models, we have found regions of the parameter space where the inflationary predictions agree with the most recent observations performed by the Planck, BICEP and Keck collaborations at the $2\sigma$ level. We have found that in order to enter the $1\sigma$ region it is necessary (and sufficient) to have a finite value of the Barbero-Immirzi parameter (a sizable $|\beta_{0}|$) and a sizable $\tilde{\xi}$ (with sign opposite to the Barbero-Immirzi parameter). This is precisely the region of parameter space for which the potential of $\chi$ develops a plateau. It is indeed the presence of this plateau that renders the theory observationally viable. As far as $\delta_{f}$ is concerned, the results of this work are obtained considering the window $\delta_{f}\in[10^{-2},16]$. We have shown that this scenario can be compatible with observations for both trans-Planckian and sub-Planckian values of $f$.
@@ Acknowledgments
The work of A.R. was supported by the Estonian Research Council grants PRG1055, RVTT3, RVTT7 and the CoE program TK202 “Fundamental Universe”. The work of A.S. was also supported by the grant DyConn from the University of Rome Tor Vergata. This article is based upon work from COST Actions COSMIC WISPers CA21106 and CosmoVerse CA21136, supported by COST (European Cooperation in Science and Technology).
@@ Appendix A A microscopic origin of natural metric-affine inflation
This relation should also be approximately true in this variant of QCD as long as the elements of $M_{q}$ are much smaller than $f$.